OpenStudy (anonymous):

solve for the variable. reduce the answer to lowest terms -15=-60

OpenStudy (eyust707):

there is no variable to solve for

OpenStudy (anonymous):

-15x = 60 x = 60/15 x = 4

OpenStudy (anonymous):

1=4?

OpenStudy (anonymous):

-15z=-60

OpenStudy (eyust707):

good guess math dude

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-15z = 60 z = 60/15 z= 4

OpenStudy (anonymous):

haha thank you

OpenStudy (anonymous):

Then @mathdood was right the first time.

OpenStudy (eyust707):

you should learn these though danielle. The fastest way to get these done is not to come on Openstudy and ask for answers

OpenStudy (anonymous):

eyust707 is right knowledge you need to have

OpenStudy (anonymous):

It is if you just have one question @eyust707

OpenStudy (anonymous):

this is** lol sound like yoda

OpenStudy (eyust707):

yea but this is the 4th one shes asked... I was helping her and she closed it to ask another on =/

OpenStudy (anonymous):

@eyust707 Well cheaters are cheaters

OpenStudy (eyust707):

yea but even cheaters work at places like the grocery store and walmart

OpenStudy (anonymous):

Where they don't don't need to know math which is good for them

OpenStudy (eyust707):

lol

OpenStudy (eyust707):

Ill stop.. I guess I just value the learning process more than most people.

OpenStudy (anonymous):

yep